OpenStudy (akonkel):

WILL MEDAL AND FAN IF YOU HAVE THE RIGHT ANSWER :) Using the slope-intercept form of a line, find the equation of the line with slope 3/2 and y-intercept -2

Mehek (mehek14):

ok so the general equation is y= mx+b m = slope b = y-intercept

Mehek (mehek14):

so now substitute 3/2 for the slope and -2 for y-intercept

OpenStudy (akonkel):

I think. Ive nver done this befor

Mehek (mehek14):

so if you substitute it, what will you get?

OpenStudy (akonkel):

I ve dont know really how to do any of this. This is my semester introduction.

Mehek (mehek14):

ok y = mx + b if you substitute the slope and y-intercept you will get y = 3/2x + (-2)

OpenStudy (akonkel):

ok :)

OpenStudy (akonkel):

i c

Mehek (mehek14):

so if you ever get another question like that, remember the general equation y=mx+b

OpenStudy (akonkel):

ok :)